# Fast Accelerated Proximal Gradient Method with New Extrapolation Term for Multiobjective Optimization

### v2

@@
Abstract: In this paper, we propose a novel extrapolation coefficient scheme within the Nesterov framework and develop an accelerated proximal gradient algorithm. We establish that the algorithm achieves a sublinear convergence rate. The proposed scheme only requires the Lipschitz constant estimate sequence to satisfy mild initial conditions, under which a key equality property can be derived to support the convergence analysis. Numerical experiments are provided to demonstrate the effectiveness and practical performance of the proposed method.
@@ 2 Preliminary Results
In this paper, for any $n\in\mathbb{N}$, $\mathbb{R}^{n}$ denotes the $n$-dimensional Euclidean space. And +$\mathbb{R}^{n}_{+}:=\{v\in\mathbb{R}^{n}\mid v_{i}\geq 0,i=1,2,\dotsb,n\}\subseteq\mathbb{R}^{n}$ signify the non-negative orthant of $\mathbb{R}^{n}$. Besides, $\Delta^{n}:=\{\lambda\in\mathbb{R}^{n}_{+}\mid\lambda_{i}\geq 0,\sum_{i=1}^{n}\lambda_{i}=1\}$ represents the standard simplex in $\mathbb{R}^{n}$. Unless otherwise specified, all inner products in this article are taken in the Euclidean space. Subsequently, the partial orders induced by +$\mathbb{R}^{n}_{+}$ are considered, where for any $v^{1},v^{2}\in\mathbb{R}^{n}$, $v^{1}\leq v^{2}$ (alternatively, $v^{1}\geq v^{2}$) holds if $v^{2}-v^{1}\in\mathbb{R}^{n}_{+}$, and $v^{1}<v^{2}$ (alternatively, $v^{1}>v^{2}$) if $v^{2}-v^{1}\in\text{int}\,\mathbb{R}^{n}_{+}$. In case of misunderstand, we define the order ($\preceq$($\prec$)⪯ ( ≺ ) in $\mathbb{R}^{n}$ as
@@
[…]
@@ 3 An accelerated proximal gradient method for multiobjective optimization
$p_{L(f)}(x,y)$ $:={\arg\min}_{z\in\mathbb{R}^{n}}\varphi_{L(f)}(z;x,y)$ (6)
$\theta_{L(f)}(x,y)$ $:={\min}_{z\in\mathbb{R}^{n}}\varphi_{L(f)}(z;x,y).$
@@
0:  Take initial point $x_{-1}=x_{0}\in\text{dom}F$, $y_{0}=x_{0}$, $\varepsilon>0$, $\gamma_{0}\in R_{++}$. Choose parameters $\eta=\frac{(k+\alpha-2)^{2}}{(k+\alpha-1)(k+\alpha-3)}$, $0<\frac{\alpha-2}{\alpha-3}s_{0}<\frac{1}{L(f)}$,if $\alpha>3$, $0<s_{0}<\frac{1}{L(f)}$, if $\alpha=3$, maximum iteration number $K$ and a vector of dimension $m$ with all components equal to one: $e_{m}$.
@@
$R:=\sup_{F^{*}\in F(X^{*}\cap\mathcal{L}_{F}(F(x_{0}))}$ $\inf_{z\in F^{-1}(\{F^{*}\})}(4\parallel 2x_{0}-z\parallel^{2}+\left\|x^{1}-z\right\|^{2})<\infty.$
4 Convergence Rate
@@
$\sigma_{k}(z)$ $:=\min_{i=1,\dotsb,m}[F_{i}(x_{k})-F_{i}(z)]$ (11)
$\rho_{p}(z)$ $:=\left\|(k+\alpha-2)x^{p}-(k+\alpha-4)x^{p-1}-z\right\|^{2}.$
@@
$-\parallel a-b\parallel^{2}+2\langle b-a,b-c\rangle=-\parallel a-c\parallel^{2}+\parallel b-c\parallel^{2}.$
@@
$-\sigma_{k+1}\geq\frac{1}{2s_{k+1}}[2\langle x_{k+1}-y_{k},y_{k}-z\rangle+\|x_{k+1}-y_{k}\|^{2}]$

### v3

@@
Abstract: In this paper, we propose a novel extrapolation coefficient scheme within a new extrapolation term and develop an accelerated proximal gradient algorithm. We establish that the algorithm achieves a sublinear convergence rate. The proposed scheme only requires the Lipschitz constant estimate sequence to satisfy mild initial conditions, under which a key equality property can be derived to support the convergence analysis. Numerical experiments are provided to demonstrate the effectiveness and practical performance of the proposed method.
@@ 1 Introduction
If we remove the smoothness assumption on $f_{i}$ and consider the case where both $f_{i}$ and $g_{i}$ are nonsmooth functions but cannot be merged into a single functional form, for example, $f_{i}(x)=\max(x,0)$ and $g_{i}(x)=\parallel x\parallel_{1}$, then traditional subdifferential based methods, such as Mäkelä et al.’s proximal bundle method [11, 12, 13, 17] and the subgradient method, become computationally demanding. This is mainly because the subdifferential is a set-valued mapping rather than a single-valued gradient, making its computation costly and intricate. To address this challenge, Gebken et al. [15] proposed a subgradient descent algorithm for nonsmooth multi-objective optimization, which combines the descent direction of [16] with an approximation based on Goldstein’s $\epsilon$-subdifferential [18].
However, such methods still suffer from the high computational burden associated with evaluating subdifferentials. This motivates a shift toward smoothing techniques [19], that is, introducing a smoothing framework to circumvent the complexity of subdifferential computations. By integrating these smoothing strategies with acceleration mechanisms, one can achieve improved convergence rates while maintaining theoretical rigor.
@@ 2 Preliminary Results
In this paper, for any $\in\mathbb{N}, n\mathbb{R}^{n}$ denotes the $n$-dimensional Euclidean space. And $\mathbb{R}^{n}_{+}:=\{v\in\mathbb{R}^{n}\mid v_{i}\geq 0,i=1,2,\dotsb,n\}\subseteq\mathbb{R}^{n}$ signify the non-negative orthant of $\mathbb{R}^{n}$. Besides, $\Delta^{n}:=\{\lambda\in\mathbb{R}^{n}_{+}\mid\lambda_{i}\geq 0,\sum_{i=1}^{n}\lambda_{i}=1\}$ represents the standard simplex in $\mathbb{R}^{n}$. Unless otherwise specified, all inner products in this article are taken in the Euclidean space. Subsequently, the partial orders induced by $\mathbb{R}^{n}_{+}$ are considered, where for any $v^{1},v^{2}\in\mathbb{R}^{n}$, $v^{1}\leq v^{2}$ (alternatively, $v^{1}\geq v^{2}$) holds if $v^{2}-v^{1}\in\mathbb{R}^{n}_{+}$, and $v^{1}<v^{2}$ (alternatively, $v^{1}>v^{2}$) if $v^{2}-v^{1}\in\text{int}\,\mathbb{R}^{n}_{+}$. In case of misunderstand, we define the order $\preceq(\prec)$ in $\mathbb{R}^{n}$ as
@@
where $\parallel\cdot\parallel$ stands for standard Euclidean norm unless specified otherwise.
@@
In the context of this study, we introduce an algorithm utilizing the smoothing function delineated in [19]. This smoothing function serves the purpose of approximating the nonsmooth convex function $f$ by a set of smooth convex functions, thereby facilitating the application of gradient-based optimization techniques.
Definition 2.1.
For convex function $f$ in (18), we call $\tilde{f}:\mathbb{R}^{n}\times\mathbb{R}_{+}\to\mathbb{R}$ a smoothing function of $f$, if $\tilde{f}$ satisfies the following conditions:
(i) for any fixed $\mu>0$$\tilde{f}(\cdot,\mu)$ is continuously differentiable on $\mathbb{R}^{n}$;
(ii) $\lim_{z\to x,\mu\downarrow 0}\tilde{f}(z,\mu)=f(x),\forall x\in\mathbb{R}^{n}$;
(iii) (gradient consistence) $\{\lim_{z\to x,\mu\downarrow 0}\tilde{f}(z,\mu)\}\subseteq\partial f(x),\forall x\in\mathbb{R}^{n}$ ;
(iv) for any fixed $\mu>0$, $\tilde{f}(z,\mu)$ is convex on $\mathbb{R}^{n}$;
(v) there exists a $k>0$ such that
$|\tilde{f}(x,\mu_{2})-\tilde{f}(x,\mu_{1})|\leq\kappa|\mu_{1}-\mu_{2}|,\forall x\in\mathbb{R}^{n},\mu_{1},\mu_{2}\in\mathbb{R}_{++};$
(vi) there exists an $L>0$ such that $\nabla_{x}\tilde{f}(\cdot,\mu)$ is Lipschitz continuous on $\mathbb{R}^{n}$ with factor $L\mu^{-1}$ for any fixed $\mu\in\mathbb{R}_{++}$.
Combining properties (ii) and (v) in Definition (2.1), we have
$|\tilde{f}(x,\mu)-f(x)|\leq\kappa\mu,\forall x\in\mathbb{R}^{n},\mu\in\mathbb{R}_{++}.$
@@ 3 An accelerated proximal gradient method for multiobjective optimization
$p_{L(f)}(x,y)$ $={\arg\min}_{z\in\mathbb{R}^{n}}\varphi_{L(f)}(z;x,y)$ (6)
$\theta_{L(f)}(x,y)$ $={\min}_{z\in\mathbb{R}^{n}}\varphi_{L(f)}(z;x,y).$
@@
0: Take initial point $x_{-1}=x_{0}\in\text{dom}F$, $y_{0}=x_{0}$, $\varepsilon>0$, $\gamma_{0}\in R_{++}$. Choose parameters $\eta=\frac{(k+\alpha-2)^{2}}{(k+\alpha-1)(k+\alpha-3)}$, $0<\frac{\alpha-2}{\alpha-3}s_{0}<\frac{1}{L(f)}$,if $\alpha>3$, $0<s_{0}<\frac{1}{L(f)}$, if $\alpha=3$, maximum iteration number $K$.
@@
$R=\sup_{F^{*}\in F(X^{*}\cap\mathcal{L}_{F}(F(x_{0}))}$ $\inf_{z\in F^{-1}(\{F^{*}\})}(4\parallel 2x_{0}-z\parallel^{2}+\left\|x^{1}-z\right\|^{2})<\infty.$
4 Convergence Rate of the Algorithm 1
@@
$\sigma_{k}(z)$ $=\min_{i=1,\dotsb,m}[F_{i}(x_{k})-F_{i}(z)]$ (11)
$\rho_{p}(z)$ $=\left\|(k+\alpha-2)x^{p}-(k+\alpha-4)x^{p-1}-z\right\|^{2}.$
@@
$\parallel a-b\parallel^{2}+2\langle b-a,b-c\rangle=-\parallel a-c\parallel^{2}+\parallel b-c\parallel^{2}.$ (15)
@@
$\sigma_{k+1}\geq\frac{1}{2s_{k+1}}[2\langle x_{k+1}-y_{k},y_{k}-z\rangle+\|x_{k+1}-y_{k}\|^{2}]$
@@
5 An accelerated proximal gradient method for nonsmooth multiobjective optimization
This section mainly discuss the following nonsmooth composite problem :
$(NCMOP)\quad\min_{x\in\mathbb{R}^{n}}F(x)\equiv(f_{1}(x)+g_{1}(x),\dotsb,f_{m}(x)+g_{m}(x))^{T},$ (18)
where $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ and $g:\mathbb{R}^{n}\to\mathbb{R},\forall i=1,\dotsb,m$ are both convex nonsmooth functions. And we will extend the algorithm in Section 3 to this case with some modifies. Surprisingly, we find that, if we choose the smoothing parameter $\mu_{k}$ and the approximate series $\{L_{k}\}$ under some assumptions, we could get some beautiful results.
Drawing inspiration from the achievements reported in [9], we incorporate extrapolation techniques with parameters $\beta_{k}=\frac{k-1}{k+\alpha-1}$, where $\alpha>3$. Choosing the smoothing function $\tilde{c}$ as defined in Definition (2.1), we formulate an accelerated proximal gradient algorithm to solve the multiobjective optimization problem denoted as (18). The algorithm achieves a faster convergence rate while also gain the sequential convergence.
Subsequently, we present the methodology employed to address the optimization problem denoted as (18). Similar to the exposition in [4], a subproblem is delineated and resolved in each iteration. Using the descent lemma, the proposed approach tackles the ensuing subproblem for prescribed values of $x\in\text{dom}(F)$, $y\in\mathbb{R}^{n}$, and $\ell\geq L$:
$\min_{z\in\mathbb{R}^{n}}\varphi_{\ell}(z;x,y,\mu),$ (19)
where
$\varphi_{l}(z;x,y,\mu):=\max_{i=1,\dotsb,m}\left[\left\langle\nabla\tilde{f}_{i}(y,\mu),z-y\right\rangle+g_{i}(z)+\tilde{f}_{i}(y,\mu)-\tilde{F}_{i}(x,\mu)\right]+\frac{\ell}{2}\left\|z-y\right\|^{2}.\\$ (20)
Since $g_{i}$ is convex for all $i=1,\dotsb,m,z\mapsto\varphi_{\ell}(z;x,y,\mu)$ is strongly convex.Thus,the subproblem (19) has a unique optimal solution $p_{\ell}(x,y,\mu)$ and attain the optimal function value $\theta_{\ell}(x,y,\mu)$,i.e.,
$p_{\ell}(x,y,\mu):=\arg\min_{z\in\mathbb{R}^{n}}\varphi_{\ell}(z,x,y,\mu)\ \text{and}\ \theta_{\ell}(x,y,\mu):=\min_{z\in\mathbb{R}^{n}}\varphi_{\ell}(z,x,y,\mu).$ (21)
Furthermore, the optimality condition associated with the optimization problem denoted as (19) implies that, for all $x\in\text{dom}\,F$ and $y\in\mathbb{R}^{n}$, there exists $\eta(x,y,\mu)\in\partial g(p_{\ell}(x,y,\mu))$ and a Lagrange multiplier $\lambda(x,y)\in\mathbb{R}^{m}$ such that
$\sum_{i=1}^{m}\lambda_{i}(x,y)[\nabla\tilde{f}_{i}(y,\mu)+\eta_{i}(x,y,\mu)]=-\ell[p_{\ell}(x,y)-y]$ (22)
$\lambda(x,y)\in\Delta^{m},\quad\lambda_{j}(x,y)=0\quad\forall j\notin\mathcal{I}(x,y),$ (23)
where $\Delta^{m}$ denotes the standard simplex and
$\mathcal{I}(x,y):=\arg\max_{i=1,\dotsb,m}[\left\langle\nabla\tilde{f}_{i}(y,\mu),p_{\ell}(x,y,\mu)-y\right\rangle+g_{i}(p_{\ell}(x,y,\mu))+\tilde{f}_{i}(y,\mu)-\tilde{F}_{i}(x,\mu)].$ (24)
For easy of reference and corresponding to its structure, we call the proposed algorithm the smoothing accelerated proximal gradient method with extrapolation term for nonsmooth multiobjective optimization(SAPGM) in this paper.The algorithm is in the following form.
Algorithm 2 The Smoothing Accelerated Proximal Gradient Method with Extrapolation term for Non-smooth Multi-objective Optimization
0: Take initial point $x_{-1}=x_{0}\in\text{dom}F$, $y_{0}=x_{0}$, $\varepsilon>0$, $\mu_{0}\in R_{++}$, $s_{0}\in R_{++}$. Choose parameters $\alpha>3$. Assume $0<\frac{\alpha-2}{\alpha-3}s_{0}\mu_{0}<\frac{1}{L(f)}$. Set maximum iteration number as $K$.
for $k=0,1,\dotsb,K$ do
2:  Compute $y_{k}=x_{k}+\frac{k+\alpha-4}{k+\alpha-1}(x_{k}-x_{k-1})$
  Compute ${x}_{k+1}=p_{L_{k}}(x_{k},y_{k},\mu_{k}),$ where $L_{k}=(s_{k}\mu_{k})^{-1}$.
4:  if $\left\|x_{k}-x_{k+1}\right\|<\varepsilon$ and $\mu_{k}<\epsilon$ then
   return $x_{k+1}$
6:  end if
  Update $\mu_{k+1}=\frac{(k+\alpha-2)\mu_{k}}{k+\alpha-1};\ s_{k+1}=\frac{(k+\alpha-2)s_{k}}{k+\alpha-3}.$
8:end for
[…]
6 Convergence rate of the algorithm 2
This section shows that SAPGM has different convergence rates with different $\sigma$ under the Assumption (3.3). For the convenience of the complexity analysis, we use some functions defined in [4]. For $k\geq 0$,for all $z\in\mathbb{R}^{n}$, let $\tilde{\sigma}_{k}:\mathbb{R}^{n}\to\mathbb{R}\cup\{-\infty\}$ be defined by
$\tilde{\sigma}_{k}(z)$ $=\min_{i=1,\dotsb,m}[\tilde{F}_{i}(x_{k},\mu_{k})-F_{i}(z)]+\kappa\mu_{k}.$ (25)
$\tilde{\rho}_{k}(z)$ $=\left\|(k+\alpha-2)x^{k}-(k+\alpha-4)x^{k-1}-z\right\|^{2}.$
Proposition 6.1.
Let $\tilde{\sigma}_{k}(z)$ define as (25), we have
$\tilde{\sigma}_{k+1}(z)\leq\tilde{\sigma}_{k}(z)-\frac{L_{k+1}}{2}[2\langle x_{k+1}-y_{k},y_{k}-x_{k}\rangle-\parallel x_{k+1}-y_{k}\parallel^{2}],$ (26)
and
$\tilde{\sigma}_{k+1}(z)\leq$ $\frac{L_{k+1}}{2}[2\langle y_{k}-x_{k+1},y_{k}-z\rangle-\|x_{k+1}-y_{k}\|^{2}]+2\kappa\mu_{k+1}.$ (27)
Proof.
Recall that there exists $\eta(x_{k},y_{k},\mu_{k+1})\in\partial g(x_{k+1})$ and Lagrange multiplier $\lambda_{i}(x_{k},y_{k})\in R^{m}$ that satisfy the KKT condition (22,23) for the subproblem (19), we have
$\tilde{\sigma}_{k+1}(z)$ $=\min_{i=1,\dotsb,m}[\tilde{F}_{i}(x_{k+1},\mu_{k+1})-F_{i}(z)]+\kappa\mu_{k+1}$
$\leq\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\tilde{F}_{i}(x_{k+1},\mu_{k+1})-F_{i}(z)+\kappa\mu_{k+1}].$
Basic from the descent lemma,
$\tilde{\sigma}_{k+1}(z)$ $\leq\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\tilde{F}_{i}(x_{k+1},\mu_{k+1})-F_{i}(z)+\kappa\mu_{k+1}]$
$=\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\tilde{F}_{i}(x_{k+1},\mu_{k+1})-\tilde{F}_{i}(z,\mu_{k+1})+\tilde{F}_{i}(z,\mu_{k+1})-F_{i}(z)+\kappa\mu_{k+1}]$
$\leq\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\tilde{F}_{i}(x_{k+1},\mu_{k+1})-\tilde{F}_{i}(z,\mu_{k+1})+2\kappa\mu_{k+1}]$
$\leq\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1}),x_{k+1}-y_{k}\rangle+g_{i}(x_{k+1})$
$\quad+\tilde{f}_{i}(y_{k},\mu_{k+1})-\tilde{F}_{i}(z,\mu_{k+1})+2\kappa\mu_{k+1}]+\frac{L_{k}}{2}\parallel x_{k+1}-y_{k}\parallel^{2}.$
Hence,the convexity of $f_{i}$ and $g_{i}$ yields
$\tilde{\sigma}_{k+1}(z)$ $\leq\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1})+\eta(x_{k},y_{k},\mu_{k+1}),x_{k+1}-z\rangle+2\kappa\mu_{k+1}]$
$\quad+\frac{L_{k}}{2}\parallel x_{k+1}-y_{k}\parallel^{2}.$
$\tilde{\sigma}_{k+1}(z)$ $\leq\frac{L_{k+1}}{2}[2\langle y_{k}-x_{k+1},y_{k}-z\rangle-\parallel x_{k+1}-y_{k}\parallel^{2}]+2\kappa\mu_{k+1},$
From the definition of $\tilde{\sigma}_{k}(z)$, we obtain
$\tilde{\sigma}_{k}(z)-\tilde{\sigma}_{k+1}(z)\geq-\max_{i=1,\dotsb,m}[\tilde{F}_{i}(x_{k+1},\mu_{k+1})-\tilde{F}_{i}(x_{k},\mu_{k})]+\kappa(\mu_{k}-\mu_{k+1})$
Basic from the descent lemma,
$\quad\tilde{\sigma}_{k}(z)-\tilde{\sigma}_{k+1}(z)$
$\geq-\max_{i=1,\dotsb,m}[\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1}),x_{k+1}-y_{k}\rangle+g_{i}(x_{k+1})+\tilde{f}_{i}(y_{k},\mu_{k+1})-\tilde{F}_{i}(x_{k},\mu_{k})]$
$\quad-\frac{L_{k+1}}{2}\parallel x_{k+1}-y_{k}\parallel^{2}+\kappa(\mu_{k}-\mu_{k+1})$
$=-\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1}),x_{k+1}-y_{k}\rangle+g_{i}(x_{k+1})+\tilde{f}_{i}(y_{k},\mu_{k+1})-\tilde{F}_{i}(x_{k},\mu_{k})]$
$\quad-\frac{L_{k+1}}{2}\parallel x_{k+1}-y_{k}\parallel^{2}+\kappa(\mu_{k}-\mu_{k+1})$
$=-\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1}),x_{k}-y_{k}\rangle+\tilde{f}_{i}(y_{k},\mu_{k+1})-\tilde{f}_{i}(x_{k},\mu_{k})]$
$\quad-\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})[\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1}),x_{k+1}-x_{k}\rangle+g_{i}(x_{k+1})-g_{i}(x_{k})]$
$\quad-\frac{L_{k+1}}{2}\parallel x_{k+1}-y_{k}\parallel^{2}+\kappa(\mu_{k}-\mu_{k+1}).$
where the first equality comes from (23),and the second one follows by taking $x_{k+1}-y_{k}=(x_{k}-y_{k})+(x_{k+1}-x_{k})$. From Definition 2.1 (v) and the convexity of $\tilde{f}_{i},g_{i}$, we show that
$\tilde{\sigma}_{k}(z)-\tilde{\sigma}_{k+1}(z)\geq$ $-\sum_{i=1}^{m}\lambda_{i}(x_{k},y_{k})\langle\nabla\tilde{f}_{i}(y_{k},\mu_{k+1})+\eta(x_{k},y_{k},\mu_{k+1}),x_{k+1}-x_{k}\rangle$
$-\frac{L_{k+1}}{2}\parallel x_{k+1}-y_{k}\parallel^{2}.$
Thus, (22) and some calculations prove that
$\tilde{\sigma}_{k}(z)-\tilde{\sigma}_{k+1}(z)\geq\frac{L_{k+1}}{2}[2\langle x_{k+1}-y_{k},y_{k}-x_{k}\rangle-\parallel x_{k+1}-y_{k}\parallel^{2}].$
So we can get two inequalities of $\tilde{\sigma}_{k}$ and $\tilde{\sigma}_{k+1}$ that we want. ∎
Theorem 6.2.
Suppose $\left\{x_{k}\right\}$ and $\left\{y^{k}\right\}$ be the sequences generated by SAPGM, for any $z\in\mathbb{R}^{n}$, it holds that
$u_{0}(x_{k+1})\leq O\left(\frac{1}{k}\right).$
Proof.
[…]
$\frac{k+\alpha-3}{k+\alpha-1}\tilde{\sigma}_{k}(z)-\tilde{\sigma}_{k+1}(z)$ (28)
$\geq$ $\frac{1}{2s_{k+1}\mu_{k+1}}\left[2\left\langle x_{k+1}-y_{k},y_{k}-\frac{k+\alpha-3}{k+\alpha-1}x_{k}-\frac{2}{k+\alpha-1}z\right\rangle+\parallel x_{k+1}-y_{k}\parallel\right]$
$\quad+\frac{4\kappa\mu_{k+1}}{k+\alpha-1}$
With equality (15), we get
$(k+\alpha-2)^{2}s_{k}\mu_{k}\tilde{\sigma}_{k}(z)-(k+\alpha-1)^{2}s_{k+1}\mu_{k+1}\tilde{\sigma}_{k+1}(z)\geq 2[\tilde{\rho}_{k+1}(z)-\tilde{\rho}_{k}(z)]+4\kappa(k+\alpha-1)\mu_{k+1}^{2}s_{k+1}.$
Applying this inequality repeatedly, we have
$s_{k+1}\mu_{k+1}(k+\alpha-1)^{2}\tilde{\sigma}_{k+1}(z)+2\tilde{\rho}_{k+1}(z)\leq s_{1}\mu_{1}W_{1}+\parallel 2x_{0}-z\parallel^{2}+4\kappa\sum_{i=1}^{k+1}(i+\alpha-1)\mu_{i}^{2}s_{i}$ (29)
And we can find that
$\mu_{i}$ $=\mu_{i-1}\frac{i+\alpha-2}{i+\alpha-1}=\mu_{0}\Pi_{j=1}^{i}\frac{j+\alpha-2}{j+\alpha-1}=\mu_{0}\frac{\alpha-1}{i+\alpha-1};$
$s_{i}$ $=s_{i-1}\frac{i+\alpha-2}{i+\alpha-3}=s_{0}\Pi_{j=1}^{i}\frac{j+\alpha-2}{j+\alpha-3}=s_{o}\frac{i+\alpha-2}{\alpha-2},$
consequently,
$4\kappa(i+\alpha-1)\mu_{i}^{2}s_{i}=4\kappa(i+\alpha-1)\frac{(\alpha-1)^{2}}{(i+\alpha-1)^{2}}\mu_{0}^{2}\frac{i+\alpha-2}{\alpha-2}s_{0}$
$=\frac{i+\alpha-2}{i+\alpha-1}4\kappa\frac{(\alpha-1)^{2}}{\alpha-2}\mu_{0}^{2}s_{0},$
we can infer that
$\sum_{i=1}^{k+1}4\kappa(i+\alpha-1)\mu_{i}^{2}s_{i}\leq O(k).$
$s_{k+1}\mu_{k+1}(k+\alpha-1)^{2}=\frac{s_{0}(k+\alpha-2)}{\alpha-2}\cdot\frac{\mu_{0}(\alpha-1)}{k+\alpha-1}\cdot(k+\alpha-1)^{2}=2(k+\alpha-1)(k+\alpha-2)\leq O(k^{2}).$
So, we get that
$\tilde{\sigma}_{k+1}(z)\leq O\left(\frac{1}{k}\right).$
which implies that
$u_{0}(x_{k+1})\leq O\left(\frac{1}{k}\right).$